# The Structure of the Majorana Clifford Group

### v2

@@ IV.2 Stabilizer Spaces & States
Finally, the projection operator satisfies
@@
In particular, for the case $r=n$ (i.e. $M$ being maximally isotropic), we can identify a single (logical) state $\lvert M,v\rangle$ satisfying (70), such that
@@
IV.3 Logical Operators
@@
One can convince oneself that $\operatorname{\mathcal{L}}$ acting on the logical Hilbert space $\operatorname{\mathcal{H}}_{\text{logic}}$ is generated by $2k$ independent Majorana strings, analogously to how there are $2n$ independent Majorana strings for the physical Hilbert space $\operatorname{\mathcal{H}}$. Furthermore, those logical operators can have odd parity iff ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ is not an element of $\operatorname{\mathcal{S}}$. This is explicitly allowed because logical operators are not necessarily observables and hence don’t have to adhere to parity superselection rules. They are therefore allowed to flip the parity of a logical state similarly to (76). From a physical perspective, this can be interpreted for example as the logical quantum states interacting with each other through the exchange of an odd number of fermions. In the existing literature this is referred to as quasiparticle poisoning [12, 15].
@@ IV.4 Stabilizers from p-Cliffords
As already implicitly shown in (64), any $[[n,0]]_{f}$ stabilizer basis can be generated by applying a single p-Clifford $\operatorname{\gamma}\in\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ to each of the standard basis strings $\operatorname{\mu}(b_{i})$ of $M_{\text{std}}$ in (65). However, things become non-trivial when considering stabilizers with $k>0$, as those might or might not contain the element ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. And since $j$ gets mapped to itself under application of any orthogonal matrix (as a direct consequence of such matrices preserving parity), applying a p-Clifford to any stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ can only result in some other stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. This is problematic e.g. if one wants to be able to prepare an arbitrary stabilizer using an encoding quantum circuit, or if one wants to uniformly sample over all possible stabilizers numerically.
There are multiple ways to approach this issue, the one we describe here is inspired by [15] and uses the addition of two unoccupied ancilla modes to the stabilizer $\operatorname{\mathcal{S}}$ to remove any potential presence of ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. One can then use (even-parity) braiding operators (58) to diagonalize the given basis of $\operatorname{\mathcal{S}}$ such that it is represented by the first $r$ rows of (65), each with two additional entries representing the ancilla appended at the end. During the process of diagonalization these ancilla modes might be occupied, but one can always choose the braiding operators such that they won’t be in the end. The composition of the inverse braiding operations (with additional phase changes) hence provides an encoding p-Clifford for $\operatorname{\mathcal{S}}$ on the extended Hilbert space.
@@
    Starting with the first basis element $S\,b_{1}=b_{1}$, because of our initial requirements we can always find a set of Householder reflections $h_{b}h_{a}$ that map it to the standard basis vector $(1,1,\ldots,0)^{T}$ (see Appendix B). We therefore add them to $S$ such that they are applied first.
@@
Since Householder reflections representing the braiding operators can be found and applied to vectors in $\mathcal{O}(n)$ time (see Appendix B), the total complexity of this algorithm is $\mathcal{O}(n^{2}r)$. In [15], a variation of this procedure is encoded as a quantum circuit, which uses the generating set of braiding operators (see Theorem 2) to perform Gaussian elimination.
@@
It is a well-known fact that the Clifford group $\operatorname{\mathcal{C}}_{2n}$ (in the fundamental representation) forms a 3-design of the unitary group $\operatorname{U}(2^{n})$ [29].
@@
An ensemble $\operatorname{\mathcal{E}}$ of unitary operators acting on a $N$-dimensional Hilbert space $\operatorname{\mathcal{H}}$ is said to be a $t$-design iff the ensemble average replicates the unitary Haar average:
@@
Equivalently, an unitary ensemble $\operatorname{\mathcal{E}}$ is a $t$-design iff the associated $t$-th frame potential
@@
matches the $t$-th frame potential $\operatorname{\mathcal{F}}_{t}(\operatorname{U}(N))\equiv\upsilon(t,N)$ of the unitary group in $N$ dimensions. The latter can be shown to have the following relevant properties [29]:
@@
Per definition, any ensemble that is a $t$-design ($t>1$) is also a $(t-1)$-design. Additionally, if the ensemble forms a group $\mathcal{G}$, $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$ simplifies to
@@
As per Lemma 2 of [29], for any subgroup $\mathcal{G}\leq\operatorname{\mathcal{C}}_{2n}$ of the Clifford group satisfying $\mathcal{G}\geq\operatorname{\mathcal{M}}_{2n}$ (including $\operatorname{\mathcal{C}}_{2n}$ itself), the frame potential can be expressed in terms of the subgroup’s symplectic representation $G\leq\operatorname{Sp}(2n)$:
@@
where $f(S)$ denotes the number of fixed points (i.e. unit eigenvectors) of $S$ in $\mathbb{F}_{2}^{2n}$. This is in part because the absolute trace $|\operatorname{tr}(\operatorname{\mu}(a)\gamma(S))|^{2}$ of a general Clifford is either zero or equal to the number of Majorana strings it commutes with.
According to the orbit-stabilizer relation, $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$ is also equivalent to the number of orbits of $G$ on $(\mathbb{F}_{2}^{2n})^{\times(t-1)}$. $\mathcal{G}$ is therefore a 2-design iff $G$ has 2 orbits on $\mathbb{F}_{2}^{2n}$ (and is transitive on $\mathbb{F}_{2}^{2n}\setminus\{0\}$). It is also a 3-design iff $G$ has 6 orbits on $(\mathbb{F}_{2}^{2n})^{\times 2}$ (or 5 if $n=1$), and so on. Using these last statements, it is easy to show that the p-Clifford subgroup is not a 2-design on $\operatorname{\mathcal{H}}$. This is because $\operatorname{O}(2n)$ has 4 orbits on $\mathbb{F}_{2}^{2n}$, corresponding to
@@
V.1 Parity-Restricted Frame Potential
However, based on the parity superselection rules we can argue that this is not the correct way to think about $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ as a $t$-design. The frame potential we are interested in should only care about how p-Cliffords (or some other ensemble) act on a (maximal) subspace of $\operatorname{\mathcal{H}}$ with fixed parity. Without loss of generality, we choose the even-parity sector for our consideration and hence define:
@@ V.1 Parity-Restricted Frame Potential
$\operatorname{\mathcal{F}}^{+}_{t}(\overline{\operatorname{\mathcal{E}}})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{E}}).$ (84)
The same approach is however not feasible for $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, as the sub-ensemble acting on the even-parity subspace is not known a priori. Instead, one can prove the following:
@@
For any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ of the p-Clifford group (including $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ itself), the parity-restricted frame potential $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ can be expressed in terms of the subgroup’s orthogonal representation $G\leq\operatorname{O}(2n)$:
@@
A formal proof is provided in Appendix A.2, but for now the expression $(f_{+}(S)+c_{+}(S))/2$ being summed over on the right side of (85) has the intuitive interpretation of being the number of fixed points in the quotient vector space $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ associated to the equivalence relation $v\sim^{c}v^{c}=v+j$. This is because the parity-check string $\operatorname{\mu}(j)\equiv(-1)^{F}$ commutes with all elements of the even-parity subgroup $\operatorname{\mathcal{M}}^{+}_{2n}$ and must therefore be equivalent to the identity $I\equiv\operatorname{\mu}(0)$ due to Schur’s lemma. This identification together with the composition property (6) then leads to the aforementioned equivalence relation and the reason why we also have to consider $c_{+}(S)$. The additional factor of 1/2 ensures that each pair $(v,v^{c})$ is only counted once, since if $v$ is a fixed/complemented point of $S$, then so is $v^{c}$.
Finally, the number of complemented points $c_{+}(S)$ is either 0 or equal to the number of fixed points $f_{+}(S)$. One can see this from the fact that the difference of two complemented points is always a fixed point, and therefore a single complemented point suffices to construct $f_{+}(S)$ distinct such points.
Using the orbit-stabilizer relation we can therefore overall conclude that $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ for any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is equal to the number of orbits of $G\leq\operatorname{O}(2n)$ on $(\mathbb{F}_{2}^{2n,+}/\sim^{c})^{\times(t-1)}$.
@@ V.2 Closing the Circle
One can now compare the parity-restricted frame potentials of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ and $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{-}$, to see if the former does a better job at replicating $\upsilon(t,2^{n-1})$ (79). This comparison can be done exactly for small $t$ and $n$ (e.g. by evaluating the orbits), but for larger values a Monte Carlo-based numerical analysis is more suitable.
What we found (exactly for $n=2,3,4$ and numerically for $n\geq 4$) is that $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})$ does match $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}_{2n}\otimes I_{-})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n})$ precisely. Using these results we were then able to formulate an exact proof:
@@
$\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n-2})\quad\forall n\geq 2,t\geq 1$ (86)
@@
Using (81) and (85), it is sufficient to show that the effective action of $\operatorname{O}(2n)$ on $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ is equal to ($2^{2n-1}$ copies of) $\operatorname{Sp}(2n-2)$, which is done in Appendix A.3. This then implies that the distributions of fixed points $f(S)$ ($S\in\operatorname{Sp}(2n)$) and $(f_{+}(S)+c_{+}(S))/2$ ($S\in\operatorname{O}(2n)$) are the same up to an overall factor of $2^{2n-1}$, which cancels with part of the denominator in (85):
@@
So while this implies that the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is also a 3-design on a fixed-parity subspace for all $n\geq 2$, there seems to be no obvious advantage compared to the naive ensemble $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{-}$, besides the former being the more natural choice due to not requiring an explicit Hilbert space decomposition.

### v3

@@ IV.2 Stabilizer Spaces & States
For the same reasons, the projection operator satisfies
@@
In particular, for the case $r=n$ (i.e. $M$ being maximally isotropic), we can identify a single (logical) stabilizer state $\lvert M,v\rangle$ satisfying (71), such that
@@
IV.3 Logical Operators & Error Correction
@@
One can convince oneself that $\operatorname{\mathcal{L}}$ acting on the logical Hilbert space $\operatorname{\mathcal{H}}_{\text{logic}}$ is generated by $2k$ independent Majorana strings, analogously to how there are $2n$ independent Majorana strings for the physical Hilbert space $\operatorname{\mathcal{H}}$. Furthermore, those logical operators can have odd parity iff ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ is not an element of $\operatorname{\mathcal{S}}$. This is explicitly allowed because logical operators are not necessarily observables and hence don’t have to adhere to parity superselection rules. They are therefore allowed to flip the parity of a logical state similarly to (77).
The same also holds true for in the context of error correction, in which case parity-flip errors can occur due to physical system interacting with the environment through the localized exchange of an odd number of Majorana modes (typically one). In the literature this is referred to as quasiparticle poisoning [12, 15], which is thought to be the primary source of errors in systems whose fundamentally Majorana degrees of freedom are sufficiently spatially separated. Fermionic codes are therefore usually constructed such that that they can detect and correct weight-1 errors, meaning they satisfy the Knill-Laflamme condition [32]:
Theorem 4 (Knill-Laflamme).
The set $\mathcal{E}$ of correctable errors of a stabilizer $\operatorname{\mathcal{S}}$ satisfies
$\langle\psi_{1}|E_{1}^{\dagger}E_{2}|\psi_{2}\rangle=\langle\psi_{1}|\psi_{2}\rangle\cdot O_{E_{1}E_{2}},$ (78)
for all error operators $E_{1},E_{2}\in\mathcal{E}$ and logical stabilizer states $\lvert\psi_{1}\rangle,\lvert\psi_{2}\rangle\in\operatorname{\mathcal{H}}_{\text{logic}}$. $O_{E_{1}E_{2}}$ is a constant depending only on $E_{1}$ and $E_{2}$.
Roughly speaking, this statement implies that for an error $E$ to be correctable there must be a unique combination of basis elements of $\operatorname{\mathcal{S}}$ that anticommute with $E$. The resulting error syndrome can then theoretically be used to find $E$ and undo it.
However these errors can accumulate over time, and if their total weight becomes larger than that of the smallest logical operator, they are not necessarily detectable anymore. One hence defines:
Definition 12 (Code Distance & Stabilizer Code).
In the context of error correction, the code distance $\delta$ of a $[[n,k]]_{f}$ stabilizer corresponds to the weight of its smallest logical operator. This is then referred to as a $[[n,k,\delta]]_{f}$ fermionic stabilizer code.
It can be shown that any error with weight smaller than $\delta$ can be detected, and any error with weight smaller than $\lfloor(\delta+1)/2\rfloor$ can also be corrected [2].
@@ IV.4 Stabilizers from p-Cliffords
As already implicitly stated in (65), any $[[n,0]]_{f}$ stabilizer basis can be generated by applying a single p-Clifford $\operatorname{\Gamma}\in\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ to each of the standard basis strings $\operatorname{\mu}(b_{i})$ of $M_{\text{std}}$ in (66). However, things become non-trivial when considering stabilizers with $k>0$, as those might or might not contain the element ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. And since $j$ gets mapped to itself under application of any orthogonal matrix (55), applying a p-Clifford to any stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ can only result in some other stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. This is problematic e.g. if one wants to be able to prepare an arbitrary stabilizer using an encoding quantum circuit, or if one wants to uniformly sample over all possible stabilizers numerically.
There are multiple ways to approach this issue, with the one we describe here being inspired by [15], meaning it incorporates additional unoccupied ancilla modes into the stabilizer $\operatorname{\mathcal{S}}$ to remove any potential presence of ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. One can then use (even-parity) braiding operators (58) to diagonalize the given basis of $\operatorname{\mathcal{S}}$ such that it is represented by the first $r$ rows of (66), each with two additional entries representing the ancilla appended at the end. During the process of diagonalization these ancilla modes might be occupied, but one can always choose the braiding operators such that they won’t be in the end. The reverse application of these operations (with additional phase changes, which we ignore) hence provides an encoding p-Clifford for $\operatorname{\mathcal{S}}$ on the extended Hilbert space.
@@
    Starting with the first basis element $S\,b_{1}=b_{1}$, because of our initial requirements we can always find a set of Householder reflections $h_{b}h_{a}$ that map it to the standard basis vector $(1,1,\ldots,0)^{T}$ (see Appendix E). We therefore add them to $S$ such that they are applied first.
@@
Since Householder reflections representing the braiding operators can be found and applied to vectors in $\mathcal{O}(n)$ time (see Appendix E), the total complexity of this algorithm is $\mathcal{O}(n^{2}r)$. In [15], this procedure is encoded as a quantum circuit which uses the generating set of braiding operators (see Theorem 2) to perform Gaussian elimination. Both approaches are not unique at all, as multiple distinct Cliffords can embed the same stabilizer.
@@
Another property of interest with regard to the parity-preserving Clifford operators is their behavior as a random unitary ensemble. This is a natural thing to investigate, as it is a well-known fact that the complete Clifford group $\operatorname{\mathcal{C}}_{2n}$ (in the fundamental representation) forms a 3-design of the (Haar-random) unitary group $\operatorname{U}(2^{n})$ [33]:
@@
An ensemble $\operatorname{\mathcal{E}}$ of unitary operators acting on a $N$-dimensional Hilbert space $\operatorname{\mathcal{H}}$ is said to be a $t$-design iff any ensemble average containing $t$ copies of $U(\cdot)U^{\dagger}$ matches the corresponding Haar-random average:
@@
From this definition it follows that any $t$-design (with $t>1$) is necessarily also a $(t-1)$-design, which can be seen by considering $O_{t}=O_{t-1}\otimes I$.
However, this characterization is hard to work with due to the presence of $O$, so instead we will be making use of the fact that an unitary ensemble $\operatorname{\mathcal{E}}$ is a $t$-design iff the associated $t$-th frame potential
@@
matches the $t$-th frame potential $\operatorname{\mathcal{F}}_{t}(\operatorname{U}(N))\equiv\upsilon(t,N)$ of the unitary group in $N$ dimensions [34]. The latter can be shown to have the following relevant properties [33, 35]:
@@
In fact, no unitary ensemble can have a frame potential smaller than $\upsilon(t,N)$, which thus provided a lower bound.
Since we are primarily interested in unitary ensembles that form a group $\mathcal{G}$ under multiplication, we can use that property to simplify $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$:
@@
This follows from the fact that $UV^{\dagger}$ is guaranteed to also be an element of the unitary ensemble in that case.
Furthermore, Lemma 2 of [33] states that for any subgroup $\mathcal{G}\leq\operatorname{\mathcal{C}}_{2n}$ of the Clifford group satisfying $\mathcal{G}\geq\operatorname{\mathcal{M}}_{2n}$, the frame potential can be expressed in terms of the subgroup’s symplectic representation $G\leq\operatorname{Sp}(2n)$:
@@
where $f(S)$ denotes the number of fixed points (i.e. unit eigenvectors) of $S$ in $\mathbb{F}_{2}^{2n}$. In the language of Majorana strings, this is in part because the absolute trace $|\operatorname{tr}U|^{2}|\equiv\operatorname{tr}(\operatorname{\mu}(a)\gamma(S))|^{2}$ of a general Clifford is either zero or equal to the number of strings it commutes with.
To deduce from (83) the frame potential of a given subgroup $\mathcal{G}$, one can employ computational methods or – in the case of small $t$ – the orbit-stabilizer relation [33]. It implies that $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$ is equivalent to the number of orbits of $G$ acting on $(\mathbb{F}_{2}^{2n})^{\times(t-1)}$, an orbit being a subset of elements that are permuted among each other under the group action. For $t=1,2,3$ this imposes the following requirements for $\mathcal{G}$ to be a $t$-design:
  • Any subgroup $\mathcal{G}$ is trivially a 1-design as $\operatorname{\mathcal{F}}_{t}(\mathcal{G})=1=\upsilon(1,N)$ for all $N$.
  • $\mathcal{G}$ is a 2-design iff $G$ has 2 orbits on $\mathbb{F}_{2}^{2n}$, i.e. it is transitive on $\mathbb{F}_{2}^{2n}\setminus\{0\}$ since $0\in\mathbb{F}_{2}^{2n}$ is always a fixed point.
  • $\mathcal{G}$ is a 3-design iff $G$ has 6 orbits on $(\mathbb{F}_{2}^{2n})^{\times 2}$, or 5 orbits if $n=1$ ($N=2$) due to (81),
With this once can deduce that $\operatorname{\mathcal{C}}_{2n}$ is indeed a 3-design on $\operatorname{\mathcal{H}}$, but $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}\leq\operatorname{\mathcal{C}}_{2n}$ is not, as $\operatorname{O}(2n)\leq\operatorname{Sp}(2n)$ has 4 orbits on $\mathbb{F}_{2}^{2n}$:
@@
This is a direct consequence of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ preserving the parity of Majorana strings, and should therefore not come as a surprise.
V.1 Fixing the Parity
However, based on the parity superselection rules we can argue that this is not the correct way to think about $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ as a $t$-design. The frame potential we are interested in should only care about how p-Cliffords (or unitaries of some other ensemble commuting with $(-1)^{F}$) act on a fixed-parity sector of $\operatorname{\mathcal{H}}$. Without loss of generality, we choose the even-parity sector for our considerations, and hence define:
@@ V.1 Fixing the Parity
$\operatorname{\mathcal{F}}^{+}_{t}(\overline{\operatorname{\mathcal{E}}})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{E}}),$ (86)
which confirms that the above definition is reasonable.
The same approach is however not feasible for $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, as the sub-ensemble acting on the even-parity subspace is not known a priori. Instead, we prove the following:
@@
For any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ of the p-Clifford group, the parity-restricted frame potential $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ can be expressed in terms of the subgroup’s orthogonal representation $G\leq\operatorname{O}(2n)$:
@@
The formal proof is provided in Appendix D.2, but for now the expression $(f_{+}(S)+c_{+}(S))/2$ being summed over on the right side of (87) has the intuitive interpretation of corresponding to the number of fixed points in the quotient vector space $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ associated to the equivalence relation $v\sim^{c}v^{c}=v+j$. This relation makes sense because in the even-parity sector of $\operatorname{\mathcal{H}}$ one has
$\operatorname{\mu}(v)\lvert\psi\rangle=\operatorname{\mu}(v)(-1)^{F}\lvert\psi\rangle=\pm\operatorname{\mu}(v^{c})\lvert\psi\rangle,$ (88)
meaning that $\operatorname{\mu}(v)$ and $\operatorname{\mu}(v^{c})$ act in the same way, up to a sign. It therefore doesn’t matter if $S$ maps $v$ to itself or its complement, which is the reason why one has to also consider $c_{+}(S)$. The additional factor of 1/2 ensures that each pair $(v,v^{c})$ is only counted once, since if $v$ is a fixed/complemented point of $S$, then so is $v^{c}$.
While the number of even-parity fixed points $f_{+}(S)$ for a given orthogonal matrix $S\in\operatorname{O}(2n)$ is not known a priori, it is easy to see that the number of even-parity complemented points $c_{+}(S)$ must be either 0 or equal to $f_{+}(S)$. One can see this from the fact that the difference of two complemented points is always a fixed point, and therefore a single complemented point suffices to construct $(f_{+}(S)-1)$ other such points.
Again using the orbit-stabilizer relation, one might conclude that $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ for any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is equal to the number of orbits of $G\leq\operatorname{O}(2n)$ on $(\mathbb{F}_{2}^{2n,+}/\sim^{c})^{\times(t-1)}$. And while this is true, things simplify significantly by closer analyzing how $\operatorname{O}(2n)$ (and its subgroups) act on this quotient space.
@@ V.2 Closing the Circle
Comparing the parity-restricted frame potentials of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ and $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{-}$ , we found (exactly for $n=2,3,4$ and approximately for $n\geq 4$) that $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})$ (using (87)) matches $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}_{2n}\otimes I_{-})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n})$ (using (83)) precisely. This lead us to formulate an exact proof:
@@
$\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n-2})\quad\forall n\geq 2,t\geq 1.$ (89)
@@
Using (83) and (87), it is sufficient to show that the effective action of $\operatorname{O}(2n)$ on $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ is equal to ($2^{2n-1}$ copies of) $\operatorname{Sp}(2n-2)$ acting on $\mathbb{F}_{2}^{2n-2}$. The proof can be found in Appendix D.3. This then implies that the distribution of $f(S)$ ($S\in\operatorname{Sp}(2n-2)$) and $(f_{+}(S)+c_{+}(S))/2$ ($S\in\operatorname{O}(2n)$) is the same up to an overall multiplicity of $2^{2n-1}$, which then cancels with part of the denominator in (87), since (41) and (53) imply that
@@
In other words, the subgroup $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ of parity-preserving Cliffords behaves like $2^{2n-1}$ copies of the ordinary Clifford group $\operatorname{\mathcal{C}}_{2n-2}$ when restricted to acting on Majorana strings with even parity and endowed with the identification $\operatorname{\mu}(v)\sim\operatorname{\mu}(v^{c})$ (up to phases) due to (88). Each copy acts the same on the even-parity strings and must therefore act differently on a single linearly-independent odd-parity string necessary to generate all of $\operatorname{\mathcal{M}}_{2n}$. The specific number of copies then arises from there being $2^{2n-1}=|\mathbb{F}_{2}^{-,2n}|$ different ways to map an odd-parity string to another odd-parity string (ignoring phases). Similar statements can probably also be made for subgroups of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, but are not of interest to us here.
Overall, while this implies that the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is indeed a 3-design on the fixed-parity subspaces of $\operatorname{\mathcal{H}}$ for all $n\geq 2$, in terms of practical applications there seems to be no obvious advantage compared to the naive ensemble $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{\pm}$, besides the former being the more natural choice due to not requiring an explicit Hilbert space decomposition. Hence there is also no advantage to simulating these ensembles on a quantum computer employing qubits degrees of freedom. Nevertheless, the recursive relationship between operators with and without parity-preservation seems to provide valuable insights into the structure of the Majorana Clifford group.